OpenStudy (anonymous):

I NEED HELPPP!! Tom has taken out a loan for college. He started paying off the loan with a first payment of $200. Each month he pays, he wants to pay back 1.2 times the amount he paid the month before. Explain to Tom how to represent his first 30 payments in sigma notation. Then explain how to find the sum of his first 30 payments.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero please help it's really important

OpenStudy (anonymous):

@phi can you please help?

OpenStudy (anonymous):

@blues

OpenStudy (anonymous):

@SolomonZelman @mathmate @Awesome781 @xGuardians @e.mccormick

OpenStudy (anonymous):

@mathmate literally anything would help <3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the first payment?

OpenStudy (anonymous):

$200

jimthompson5910 (jim_thompson5910):

What's the next payment

OpenStudy (phi):

and the 2nd payment ?

OpenStudy (anonymous):

idk.. the second payment needs to be 1.2 times more than the first payment, soooo 240?

jimthompson5910 (jim_thompson5910):

and the next?

OpenStudy (anonymous):

288 but idk what sigma notation is :(

jimthompson5910 (jim_thompson5910):

in general, do you know how to find the payment for any given month?

OpenStudy (anonymous):

ummm I dont think so

jimthompson5910 (jim_thompson5910):

Payment 1: 200 Payment 2: 240 = 200*1.2 Payment 3: 288 = 240*1.2 etc etc

jimthompson5910 (jim_thompson5910):

notice how payment 3 can be written as 288 = 240*1.2 = (200*1.2)*1.2

jimthompson5910 (jim_thompson5910):

Then (200*1.2)*1.2 = 200*(1.2*1.2) = 200*(1.2)^2

jimthompson5910 (jim_thompson5910):

What's payment 4?

OpenStudy (anonymous):

about 345. are you sure thats sigma notation?

jimthompson5910 (jim_thompson5910):

we'll get there

jimthompson5910 (jim_thompson5910):

288*1.2 = 345.6 We can write that as 345.6 = 288*1.2 = ((200*1.2)*1.2)*1.2 = 200*(1.2)^3 starting to see a pattern?

OpenStudy (phi):

jim is explaining how to find each payment. after you know that, we use a short-cut notation to write the answer that will be the sigma notation. But you have to know what the series is to write it down. Once you know it, the sigma part is relatively easy

OpenStudy (anonymous):

oh okay. I think I get it now. Thank you all so much. I have one more problem can you guys please help? It's the last one

OpenStudy (anonymous):

Bobby has been planting blackberries in a garden in his backyard. Bobby started with blackberries in one square foot of garden modeled by the function s(x) = 30. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blackberry plants to produce more by a rate of a(x) = 1.3^(x-1). Explain to Bobby how to create an equation to predict the number of blackberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blackberries he will grow with 20 square feet.

OpenStudy (anonymous):

I know its long but I dont think its that complicated. please I really need help

jimthompson5910 (jim_thompson5910):

before we move on, what's the nth term? What's the general formula to get any payment you want?

OpenStudy (anonymous):

uh oh idk :o

OpenStudy (anonymous):

like 200 * 1.2^2

jimthompson5910 (jim_thompson5910):

that's a specific payment I want the general payment for any month

OpenStudy (anonymous):

I have no clue

jimthompson5910 (jim_thompson5910):

First payment: 200 = 200*(1.2)^0 Second payment: 240 = 200*(1.2)^1 Third payment: 288 = 200*(1.2)^2 Fourth payment: 345.6 = 200*(1.2)^3 See the pattern?

OpenStudy (anonymous):

yeah, but whats the general payment?

jimthompson5910 (jim_thompson5910):

what's the pattern? or the basic template?

OpenStudy (anonymous):

please just tell me lol Im like helen keller

OpenStudy (anonymous):

wait so can you tell me how to represent these payments in sigma notation? I get everything else

jimthompson5910 (jim_thompson5910):

What's the same each time? What's different? look at the list again First payment: 200 = 200*(1.2)^0 Second payment: 240 = 200*(1.2)^1 Third payment: 288 = 200*(1.2)^2 Fourth payment: 345.6 = 200*(1.2)^3

OpenStudy (anonymous):

youre always multiplying by 200 and 1.2 and the exponent is different every time

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so can you make a guess where the variable goes?

jimthompson5910 (jim_thompson5910):

keep in mind that variable means "something that changes" (vary ---> vari-able, vary = change)

OpenStudy (anonymous):

the exponent?

OpenStudy (anonymous):

it's the only thing changing

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so for the first payment 200 = 200*(1.2)^0 if n = 1 represents the first payment, then n-1 must be the exponent (to get 0)

jimthompson5910 (jim_thompson5910):

That's why the general nth term is \[\large 200(1.2)^{n-1}\]

OpenStudy (anonymous):

oh okay. I get it. Can you tell me how to represent it in sigma notation now?

jimthompson5910 (jim_thompson5910):

well you use the symbol \(\Large \sigma\) this is the greek upper case letter sigma

jimthompson5910 (jim_thompson5910):

Below the sigma, you tell where to start You start at n = 1

jimthompson5910 (jim_thompson5910):

Above the sigma, you tell where to stop. We stop at n = 30 (because we have 30 payments)

jimthompson5910 (jim_thompson5910):

Off to the right, you write what you want to sum up, which is \[\large 200(1.2)^{n-1}\]

jimthompson5910 (jim_thompson5910):

Put this all together to get \[\large \sum_{n=1}^{30} 200(1.2)^{n-1}\]

jimthompson5910 (jim_thompson5910):

That's the shorthand way to say "add up the payments from n = 1 to n = 30 from the nth term formula \(\Large 200(1.2)^{n-1}\) "

OpenStudy (anonymous):

okay, I completely understand now. Thank you so much. Do you have any idea on that question I posted before?

jimthompson5910 (jim_thompson5910):

oops I meant to write \(\Large \Sigma\) and NOT \(\Large \sigma\)

OpenStudy (anonymous):

would the series be convergent or divergent? im stuck on that part..

jimthompson5910 (jim_thompson5910):

@gennybaby99 notice how r = 1.2 since |r| < 1 is NOT true, this means the series does NOT converge The series diverges. If you keep going forever, the sum grows and grows. It doesn't approach a single fixed number

OpenStudy (anonymous):

thank you soo much